Find the area of the walkway (24" wide) that will be installed around a circular Jacuzzi with diameter of 6'.
Answer by Cromlix(4381)

Due to the width of the path the diameter
increases to 6' + 2' + 2' = 10'
Area of whole area = Pi * 5^2 = 78.54ft^2
area of jacuzzi = Pi * 3^2 = 28.27ft^2
78.54 - 28.27 = 50.27ft^2
This is the area of the pathway.
Hope this helps.
:-)